OpenStudy (anonymous):

(7 - √54) - (5 + √24)

OpenStudy (cathyangs):

You combine like terms...so this would be 2-√54-√24. However, you probably have to simplify the radicals to: 2-3√6-2√6=2-√6

OpenStudy (anonymous):

\[(7 - 3\sqrt{6}) - (5 + 2\sqrt{6})\]\[7 - 3\sqrt{6} - 5 - 2\sqrt{6} = 2 - 5\sqrt{6}\]

OpenStudy (cathyangs):

oops...it's 2-5√6 :d